# Pranav Gilda    Prakhar Dungarwal    Ariel Thongkham    Elena T. Ajayi    Shaifali Choudhary    Teresa Mondria Terol    Christine Lam    Jessica Pardim Araujo    Megan McFadyen-Mungall    Larry S. Liebovitch    Peter T. Coleman    Harry West    Kate Sieck    Scott Carter

### v1

@@ Abstract
We used machine learning and artificial intelligence: 1) to measure levels of peace in countries from news and social media and 2) to develop on-line tools that promote peace by helping users better understand their own media diet.
For news media, we used neural networks to measure levels of peace from text embeddings of on-line news sources. The model, trained on one news media dataset also showed high accuracy when used to analyze a different news dataset.
@@
To promote peace, we note that 71% of people 20-40 years old daily view most of their news through short videos on social media. Content creators of these videos are biased towards creating videos with emotional activation, making you angry to engage you, to increase clicks. We developed and tested a Chrome extension, MirrorMirror, which provides real-time feedback to YouTube viewers about the peacefulness of the media they are watching. Our long term goal is for MirrorMirror to evolve into an open-source tool for content creators, journalists, researchers, platforms, and individual users to better understand the tone of their media creation and consumption and its effects on viewers. Moving beyond simple engagement metrics, we hope to encourage more respectful, nuanced, and informative communication.
@@ I Introduction
In this paper, we now report how we extended that previous work to analyze quantitative levels of peace and social dimensions important in peace from on-line news sources and social media such as text transcripts of YouTube videos, using more sophisticated machine learning and AI methods:
@@ II-A Strategy
There are many surveys that measure levels of peace in a country conducted by international organizations, national governmental agencies, non-profit groups, and for-profit consulting companies [1]. Thus, we can use that data to tag samples of text from different countries with their level of peace to form training and testing sets for supervised learning. We also tested the accuracy of models trained on one dataset analyzing another dataset of different countries.
@@ II-B3 Training Procedure
III Results: News Data
III-A Evaluation of Neural Networks for Peace Classification
III-A1 Model Performance on Test Data
Across the three network architectures, all models achieved high performance on the held-out test set from the NOW dataset. The convolutional neural network (CNN) achieved a test accuracy of 97.24%, the fully connected feed-forward network reached 97.48%, and the revised CNN achieved 96.99%. These results indicate that the models effectively learned to distinguish linguistic patterns associated with high- and low-peace discourse within structured news text.
III-A2 Cross-Dataset Generalization
TABLE I: Neural Network Performance Across Datasets
Model NOW (%) Capstone (%)
CNN 97.24 72.81
Feed-Forward 97.48 72.47
Revised CNN 96.99 69.91
III-A3 Comparison Between Network Architectures
III-A4 Cross-Domain Generalization to YouTube Transcripts
IV Methods: YouTube Data
IV-A Strategy
The news data models were trained by tagging news data with the levels of peace from external traditional peace indices. There is no similar external peace measures to tag YouTube videos. Therefore, we based our YouTube models both on 1) feedback that we received from peace experts and journalists reviewing our previous machine learning results and 2) on social dimensions identified by 59 studies over the last 40 years as important measures of the level of peace, including studies on: integrative complexity, political framing meta-analysis, tightness/looseness, and many others [12][13][14][15]. These dimensions are:
@@ IV-A Strategy
IV-B Analytical Models for YouTube Transcripts
@@ IV-C Google GoEmotions
IV-D Large Language Models (LLMs)
@@ IV-E Human Feedback
Inter-rater reliability analysis validated the robustness of this gold standard, yielding pairwise correlations exceeding $r=0.93$ for three dimensions (News, Compassion, Order) and 100% agreement within a single-point margin. This high consensus confirms that our peace dimensions are observable and measurable constructs, distinct from random subjective noise. We aggregated human scores by averaging across all available raters for each video-dimension combination, with sample sizes ranging from N=32 to N=47 depending on missing data patterns. The resulting gold standard dataset provides mean scores, standard deviations, and distributional statistics that serve as ground truth for evaluating model performance.
TABLE II: Human Expert Annotation Statistics

### v2

@@ Abstract
We used machine learning and artificial intelligence: 1) to measure levels of peace in countries from news and social media and 2) to develop online tools that promote peace by helping users better understand their own media diet.
For news media, we used neural networks to measure levels of peace from text embeddings of online news sources. The model, trained on one news media dataset also showed high accuracy when used to analyze a different news dataset.
@@
To promote peace, we note that 71% of people 16-40 years old daily view most of their news through short videos on social media. Content creators of these videos are biased towards creating videos with emotional activation, making you angry to engage you, to increase clicks. We developed and tested a Chrome extension, MirrorMirror, which provides real-time feedback to YouTube viewers about the peacefulness of the media they are watching. Our long term goal is for MirrorMirror to evolve into an open-source tool for content creators, journalists, researchers, platforms, and individual users to better understand the tone of their media creation and consumption and its effects on viewers. Moving beyond simple engagement metrics, we hope to encourage more respectful, nuanced, and informative communication.
@@ I Introduction
In this paper, we now report how we extended that previous work to analyze quantitative levels of peace and social dimensions important in peace from online news sources and social media such as text transcripts of YouTube videos, using more sophisticated machine learning and AI methods:
@@ II-A Strategy
There are many surveys that measure levels of peace in a country conducted by international organizations, national governmental agencies, non-profit groups, and for-profit consulting companies [1]. Thus, we can use that data to tag samples of text from different countries with their level of peace to form training and testing sets for supervised learning. We also tested the accuracy of models trained on one dataset analyzing another dataset of different countries, see Fig. 1.
@@ II-B3 Training Procedure
III Methods: YouTube Data
III-A Strategy
The news data models were trained by tagging news data with the levels of peace from external traditional peace indices. There are no similar external peace measures to tag YouTube video transcripts. Therefore, we based our YouTube models both on 1) feedback that we received from peace experts and journalists reviewing our previous machine learning results and 2) on social dimensions identified by 59 studies over the last 40 years as important measures of the level of peace, including studies on: integrative complexity, political framing meta-analysis, tightness/looseness, and many others [12][13][14][15]. These dimensions are:
@@ III-A Strategy
III-B Analytical Models for YouTube Transcripts
@@ III-C Google GoEmotions
III-D Large Language Models (LLMs)
@@ III-E Human Feedback
Inter-rater reliability analysis validated the robustness of this gold standard, yielding pairwise correlations exceeding $r=0.93$ for three dimensions (News, Compassion, Order) and 100% agreement within a single-point margin. This high consensus confirms that our peace dimensions are observable and measurable constructs, distinct from random subjective noise. We aggregated human scores by averaging across all available raters for each video-dimension combination, with sample sizes ranging from N=32 to N=47 depending on missing data patterns. The resulting gold standard dataset provides mean scores, standard deviations, and distributional statistics that serve as ground truth for evaluating model performance, as shown in Table I.
TABLE I: Human Expert Annotation Statistics
@@
IV Results: News Data
IV-A Evaluation of Neural Networks for Peace Classification
IV-A1 Model Performance on Test Data
Across the three network architectures, all models achieved high performance on the held-out test set from the NOW dataset, as shown in Table II, The convolutional neural network (CNN) achieved a test accuracy of 97.24%, the fully connected feed-forward network reached 97.48%, and the revised CNN achieved 96.99%. These results indicate that the models effectively learned to distinguish linguistic patterns associated with high- and low-peace discourse within structured news text.
IV-A2 Cross-Dataset Generalization
TABLE II: Neural Network Performance Across Datasets
Model NOW (%) Capstone (%)
CNN 97.24 72.81
Feed-Forward 97.48 72.47
Revised CNN 96.99 69.91
IV-A3 Comparison Between Network Architectures
IV-A4 Cross-Domain Generalization to YouTube Transcripts